You have a 5 mM stock of MgCl2 and you want to prepare a 30 ul mixture of PCR reagents that has a final concentration of 100 uM of MgCl2. How much of the stock MgCl2 would you need to add to the PCR mixture?
Answer #1

According to the formula:

M1V1 = M2V2,

5 * 10-3 * V1 = 100uM * 30

V1 = 100*30/5 * 10-3 = 0.6 ul.

Thus a total of 0.6ul of stock solution can be used for the working reaction to make a solution of 30ul. However, this volume cannot be pipetted out so the stock solution can be diluted at least 10 times to get a working diluted solution, and then a volume of 6ul can be pipetted for the working reaction.